OpenStudy (anonymous):

are 2x=5y-3 and 15y=9-6x parallel lines

OpenStudy (anonymous):

Look : we have : y=-2/5x+3/5 y=-6/15x9/15 =>y=-2/5x * 3/5 To have parallel lines we should have : m1=m2 Just it ! So what do you think ?! Can you find the answer ? WE have : 2/5=2/5

OpenStudy (anonymous):

@ashleyarntz21